OpenStudy (anonymous):

y=-2x+1 find x and y intercepts thanks :)

sam (.sam.):

y=-2x+1 y=mx+c c is the y-intercept

OpenStudy (anonymous):

put x=0, we can get y-intercept which is 1 put y=0, we can get x-intercept which is 1/2

OpenStudy (anonymous):

so -1/-2 is x?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so is it -1/-2 or 1/2

OpenStudy (anonymous):

it is the same as (-1)/(-2)=(-1)(-1) / (-1)(-2) = 1/2

OpenStudy (anonymous):

ok thanks so the coordinates are 1/2,0 and 0,1?